OpenStudy (anonymous):

What is the general rule for the arithmetic sequence? 15, 12, 9, 6, … A. 3n – 18 B. 6(3 – n) C. 3n + 12 D. 18 – 3n

OpenStudy (mathstudent55):

Try each one for a few terms, and see which one works.

OpenStudy (anonymous):

-3

OpenStudy (aum):

15, 12, 9, 6, … What is the common difference d?

OpenStudy (anonymous):

you subtract 3 so its -3

OpenStudy (anonymous):

B

OpenStudy (aum):

Once you find d, the nth term of an arithmetic sequence is: \(\Large a_n = a_1 + (n-1)d\) Substitute for a1 and d and simplify.

OpenStudy (mathstudent55):

Let's look at A: 3n – 18 n = 1: 3 - 18 = -15 Since the first term is 15, not -15, this can't be it.

OpenStudy (mathstudent55):

Let's look at B: 6(3 - n) n = 1: 6(3 - 1) = 6 * 2 = 12. Since the first term is 15, not 12, this can't be it.

OpenStudy (mathstudent55):

Do the same for choices C and D. If with one choice, n = 1 works, then use n = 2, 3, 4, and make sure all of the n values also work.